OpenStudy (anonymous):

Find the area : Trapezoid: b(1)= 9 7/8, b(2)= 5 3/4, h= 6 2/3

OpenStudy (phi):

First, you need the formula for the area of a trapezoid (or you need to figure it out, which is possible but harder to do)

OpenStudy (anonymous):

I know the formula hun

OpenStudy (phi):

do you know the formula for the area of a trapezoid ? http://www.mathgoodies.com/lessons/vol1/area_trapezoid.html it is A= ½ * (b1 + b2) * h replace b1 and b2 with the numbers replace h with the numbers change the mixed numbers to improper fractions add the fractions b1 + b2 (you will need a common denominator... use 4) if you get this far , what do you have ?

OpenStudy (phi):

* make the common denominator 8 (not 4) add the fractions b1 + b2 (you will need a common denominator)

OpenStudy (anonymous):

ive did part of what u said and I now have A= 1/2(79/8 + 23/4) *20/3

OpenStudy (phi):

looks good so far. you want 23/4 to have a denominator of 8 so that it matches 79/8 multiply BOTH top and bottom by 2 to get 46/8 now you can add the tops

OpenStudy (anonymous):

Ok I got that so far

OpenStudy (phi):

what do you have now ?

OpenStudy (anonymous):

I added the numerators and I got 125/8

OpenStudy (phi):

yes, so the problem is now \[ \frac{1}{2} \frac{125}{8}\frac{20}{3} \] I would simplify first, and then multiply the tops, and multiply the bottoms

OpenStudy (anonymous):

Okay thanks ill let u know what I come up with in a sec hun , sorry for the wait I had to discuss somethings with my family

OpenStudy (anonymous):

hey hun I cant simplify if I do then what should I do ?